# Targeted Distillation for Sentiment Analysis

### v1

@@ Abstract
This paper presents a compact model that achieves strong sentiment analysis capabilities through targeted distillation from advanced large language models (LLMs). Our methodology decouples the distillation target into two key components: sentiment-related knowledge and task alignment. To transfer these components, we propose a two-stage distillation framework. The first stage, knowledge-driven distillation (KnowDist), transfers sentiment-related knowledge to enhance fundamental sentiment analysis capabilities. The second stage, in-context learning distillation (ICLDist), transfers task-specific prompt-following abilities to optimize task alignment. For evaluation, we introduce SentiBench, a comprehensive sentiment analysis benchmark comprising 3 task categories across 12 datasets. Experiments on this benchmark demonstrate that our model effectively balances model size and performance, showing strong competitiveness compared to existing small-scale LLMs. [^1] We will release our code, data, and model weights via Github.
@@
Yice Zhang1,3thanks:  The first two authors contribute equally to this work. , Guangyu Xie1,3∗, Jingjie Lin1,3, Jianzhu Bao1,3, Qianlong Wang1,3, Xi Zeng4, and Ruifeng Xu1,2,3thanks:  Corresponding Authors 1 Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China 3 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 4 The 30th Research Institute of China Electronics Technology Group Corporation zhangyc_hit@163.com,guangyuxie2001@gmail.com,xuruifeng@hit.edu.cn
@@ 1 Introduction
Sentiment analysis, aiming to identify and extract subjective information from user-generated content Liu (2012), has emerged as a significant research area in natural language processing, garnering widespread attention Zhang et al. (2018); Wankhade et al. (2022). Recent studies demonstrate that large language models (LLMs) exhibit remarkable capabilities and achieve state-of-the-art performance in sentiment analysis tasks Zhang et al. (2024b); Wang et al. (2024c); Šmíd et al. (2024). Despite these advancements, the practical application of LLMs faces significant challenges. Deploying these models incurs considerable computational costs, and fine-tuning them for enhanced task-specific performance demands greater computational resources.
Figure 1: The comparison of our model with other small-scale models in terms of the average performance on SentiBench ($F_{1}$-score, %).
@@
Motivated by these insights, this paper explores targeted distillation specifically for sentiment analysis. We decouple the distillation target into knowledge and alignment and propose a two-stage distillation framework. The first stage, termed knowledge-driven distillation (KnowDist), focuses on transferring fundamental sentiment analysis capabilities, thereby improving the student model’s potential performance. In KnowDist, we devise a multi-perspective prompting strategy to elicit comprehensive sentiment-related knowledge from the teacher LLM and systematically transfer this knowledge to the student model. The second stage, termed in-context learning distillation (ICLDist), transfers prompt-following capabilities in sentiment analysis to optimize the student model’s task alignment. In ICLDist, we enable the student model to follow task-specific instructions and demonstrations by mimicking the teacher LLM’s responses on few-shot samples. When constructing few-shot samples, we implement format and task diversification strategies to strengthen the generalization of ICLDist.
To facilitate a systematic evaluation, we develop SentiBench, a comprehensive sentiment analysis benchmark. Our extensive experimentation on this benchmark reveals several key findings: (1) Our approach demonstrates substantial advantages over generic distillation methods, achieving effective distillation of LLMs’ sentiment analysis capabilities. Specifically, the student model achieves a 10% improvement in the average $F_{1}$-score across various tasks, with a particularly remarkable gain of 38% in irony detection. (2) Leveraging our approach, Llama-3-1.2B outperforms Llama-3-3.2B and exhibits strong competitiveness against other small-scale models (see Figure 1). (3) Further analysis reveals the complementary nature of KnowDist and ICLDist and validates the effectiveness of each component in our approach.
@@ 2 Two-stage Distillation Framework
${\cal L}(\theta_{S})$ $=-\sum_{\bm{i},\bm{d},\bm{x},\bm{y}}\log P_{\cal M}(\bm{y}\ |\ \bm{i},\bm{d},\bm{x};\theta_{S}),$ (2)
$\hat{\theta}_{S}$ $=\underset{\theta_{S}}{\operatorname{argmin}}\ {\cal L}(\theta_{S}),$ (3)
where $\cal M$ denotes the teacher or student model, and $\theta_{T}$ and $\theta_{S}$ denote their respective parameters.
@@ 2.2 In-Context Learning Distillation
[…]
@@
Structured sentiment analysis Opener 1744 249 499 - micro_f1
Table 1: Task overview and dataset statistics in SentiBench. We perform downsampling on some datasets to ensure computational efficiency. For sampling details, please refer to Appendix B.
@@
The ICLDist process is illustrated in Figure 2. Similar to knowledge collection, we first gather a large volume of user-generated content. Next, we select sentiment classification and emotion recognition as distillation tasks and construct prompts by randomly applying our format diversification strategies. Additionally, we incorporate the task diversification strategy to generate supplementary prompts. We then collect the teacher LLM’s responses to these prompts, resulting in a task-alignment corpus. Finally, we optimize the student model on this corpus to enhance its task alignment.
@@ 3 SentiBench
Table 2: Experimental results on SentiBench ($F_{1}$-score, %). All results except T5-base are obtained through in-context learning with 4 demonstrations. Our distillation uses 1M KnowDist and 500K ICLDist samples.
@@ 4.1 Experimental Setup
Implementation Details. We employ Llama-3.1-70B-Instruct Grattafiori et al. (2024) as the teacher LLM and Llama-3.2-1.2B-Instruct [^2] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ as the student model. For distillation, we curate a large and diverse corpus of user-generated texts from IMDb Nguyen et al. (2014), Yelp [^3] https://www.yelp.com/dataset, Amazon [^4] https://nijianmo.github.io/amazon/index.html, and Twitter [^5] https://archive.org/details/twitterstream. We preprocess this corpus by removing texts that overlap with downstream datasets, eliminating duplicates using simhash, and filtering out too short or long texts. We then apply the proposed prompting methods to these user texts and obtain 1M KnowDist samples and 400K ICLDist samples. We further supplement the ICLDist corpus with 100K general task samples from the super-naturalInstruction Wang et al. (2022) dataset. Subsequently, we conduct a two-stage optimization of the student model using these two corpora. For the main experiments, we utilize the complete dataset, while for ablation studies, we leverage a subset consisting of 200K KnowDist samples and 100K ICLDist samples. The hyperparameter settings are provided in Appendix C.
After distillation, we evaluate the student model on SentiBench using in-context learning, with dataset statistics shown in Table 1. The specific prompts for each task are provided in Appendix B. During evaluation, we randomly sample 4 examples from the training set as demonstrations. To ensure generation stability, we set the temperature parameter to 0 during model inference. To mitigate the impact of randomness, we conduct each evaluation using 3 different random seeds and report the averaged results.
Baselines. We compare our approach with generic distillation methods. Specifically, we train the student model using existing instruction-following datasets, including the 52K data constructed by Taori et al. (2023) (alpaca-data), and the 2.58M data developed by Wu et al. (2024) (lamini-data). Furthermore, we evaluate a diverse set of models for reference: (1) a fine-tuned T5-base model Raffel et al. (2020); (2) models from the Llama-3 family, spanning different scales (3.2B, 8B, and 70B variants); (3) several small-scale models ranging from 1B to 3B parameters, including OPT-1.3B Zhang et al. (2022), TinyLlama-1.1B-Chat-v1.0 Zhang et al. (2024a), Phi-2-2.7B [^6] https://huggingface.co/microsoft/phi-2, Qwen-2.5-1.5B-Instruct [^7] https://qwenlm.github.io/blog/qwen2.5/, and Gemma-2-2.6B-it Team (2024); and (4) GPT-3.5888Available at https://chat.openai.com/. The specific model used is gpt-3.5-turbo-0125..
@@ 4.2 Main Results
Table 2 presents the comparison results on SentiBench. We observe that two generic distillation methods yield only marginal gains in sentiment analysis performance, with the student model showing average $F_{1}$-score improvements of 0.81% and 1.43% respectively. These limited improvements suggest that utilizing generic distillation methods to transfer sentiment analysis capabilities is challenging. In contrast, our approach, namely Know & ICLDist, significantly enhances the sentiment analysis performance of the student model. Specifically, our approach achieves an average improvement of over 10%. The most striking improvement is observed in irony detection, where the $F_{1}$-score increases dramatically from 35.80% to 73.80% - an improvement of 38.00%. These results demonstrate the effectiveness of our approach in transferring sentiment analysis capabilities from the LLM to its more efficient counterparts.
Furthermore, the experimental results in Table 2 reveal several additional insights. Firstly, within the Llama-3 family, we observe a clear positive correlation between model size and performance, with Llama-3-70B achieving the best results, surpassing GPT-3.5 and approaching the fine-tuned T5-base model. Secondly, with Llama-3-70B as the teacher LLM, our approach enables Llama-3-1.2B to achieve comparable performance to the teacher on sentiment classification and emotion recognition. Thirdly, our approach empowers Llama-3-1.2B to outperform Llama-3-3.2B. Moreover, the distilled model demonstrates strong competitive performance compared to other small-scale models and GPT-3.5, also illustrated in Figure 1. Finally, both the distilled model and other small-scale models show inferior performance on intimacy analysis and tuple extraction tasks (i.e., ASQP and SSA). These tasks require a deep understanding of social context and advanced structured extraction capabilities, presenting promising directions for future research.
@@ 4.3 Analyis of Two-stage Optimization
Our distillation framework consists of two stages: KnowDist and ICLDist. The results in Table 2 demonstrate that both stages enhance the student model’s performance on sentiment analysis tasks, and combining them yields even better results. Below, we conduct an in-depth analysis of these two stages, aiming to distinguish their respective roles and investigate how they complement each other.
@@
For the first question, we conduct fine-tuning experiments using the training samples from SentiBench. The results in Table 3 demonstrate that under the fine-tuning setting, both KnowDist and ICLDist can enhance the student model’s sentiment analysis performance. Notably, KnowDist achieves more substantial improvements, which contrasts with the in-context learning results in Table 2. These findings support our claims: KnowDist strengthens the student model’s fundamental sentiment analysis capabilities, while ICLDist optimizes task alignment. When sufficient labeled samples are available for downstream task alignment, the benefits of ICLDist’s task alignment become less significant. However, such labeled data is often scarce in real-world applications. Consequently, both KnowDist and ICLDist are essential components of our framework.
@@ 4.4 Ablation Studies
Table 5: Ablation results of KnowDist ($F_{1}$-score, %). Anl and Rw denote analyzing and rewriting respectively, and MPP stands for multi-perspective prompting.
@@
Table 6: Ablation results of ICLDist ($F_{1}$-score, %). LW, LT, and MI denote the format diversification of Label Words, Label Taxonomies, and Minimized Instructions respectively, while TD represents Task Diversification. We divide tasks in SentiBench into seen and unseen categories during distillation, where seen tasks include sentiment classification and emotion recognition, while the rest are considered unseen.
@@ 4.5 Discussions
Effect of Teacher LLMs. We experiment with different teacher LLMs in our distillation framework to analyze their impact. The results in Table 7 reveal that teacher quality significantly influences distillation effectiveness, as larger teacher LLMs generally lead to more substantial improvements. Furthermore, we make two noteworthy observations. First, even when using identical models for both teacher and student, distillation has the potential to enhance the student’s sentiment analysis performance. This result suggests the potential for leveraging distillation to achieve self-improvement in specialized domains. Second, higher teacher quality does not always correlate with better student performance, as evidenced by the FSA performance when using 8B and 70B models as teachers. This may be related to the capability gap between teachers and students, warranting further exploration in future work.
@@
Table 8: Experimental results on 5-shot MMLU (accuracy, %). Our evaluation is conducted using LM-Evaluation-Harness provided at %****␣main.tex␣Line␣575␣****https://github.com/meta-llama/llama-cookbook.

### v2

@@ Abstract
This paper explores targeted distillation methods for sentiment analysis [^1] In this paper, we adopt a broad definition of sentiment analysis, which encompasses not only traditional polarity classification but also a range of related tasks such as emotion recognition, irony detection, and stance detection., aiming to build compact and practical models that preserve strong and generalizable sentiment analysis capabilities. To this end, we conceptually decouple the distillation target into knowledge and alignment and accordingly propose a two-stage distillation framework. Moreover, we introduce SentiBench, a comprehensive and systematic sentiment analysis benchmark that covers a diverse set of tasks across 12 datasets. We evaluate a wide range of models on this benchmark. Experimental results show that our approach substantially enhances the performance of compact models across diverse sentiment analysis tasks, and the resulting models demonstrate strong generalization to unseen tasks, showcasing robust competitiveness against existing small-scale models. [^2] We release our code, data, and model weights at https://github.com/HITSZ-HLT/Sentiment-Distillation.
@@
Yice Zhang1thanks:   The first two authors contribute equally to this work. , Guangyu Xie1∗, Jingjie Lin1, Jianzhu Bao4,1, Qianlong Wang1, Xi Zeng3, and Ruifeng Xu1,2thanks:   Corresponding Authors 1 Harbin Institute of Technology, Shenzhen, China 2 Peng Cheng Laboratory, Shenzhen, China 3 The 30th Research Institute of China Electronics Technology Group Corporation 4 Nanyang Technological University, Singapore zhangyc_hit@163.com,guangyuxie2001@gmail.com,xuruifeng@hit.edu.cn
@@ 1 Introduction
Sentiment analysis, aiming to identify and extract subjective information from user-generated content Liu (2012), has emerged as a significant research area in natural language processing, garnering widespread attention Zhang et al. (2018); Wankhade et al. (2022). Recent studies demonstrate that large language models (LLMs) exhibit remarkable capabilities and achieve state-of-the-art performance in diverse sentiment analysis tasks Zhang et al. (2024b); Wang et al. (2024c); Šmíd et al. (2024). Despite these advancements, the practical application of LLMs faces significant challenges. Deploying these models incurs considerable computational costs, and fine-tuning them for enhanced task-specific performance demands greater computational resources.
Figure 1: The comparison of our distilled model with other small-scale models in terms of the average performance on SentiBench ($F_{1}$-score, %).
@@
Motivated by these insights, this paper explores targeted distillation specifically for sentiment analysis. We conceptually decouple the distillation target into knowledge and alignment and propose a two-stage distillation framework. The first stage, termed knowledge-driven distillation (KnowDist), focuses on transferring fundamental sentiment analysis capabilities, thereby improving the student model’s potential performance. In KnowDist, we devise a multi-perspective prompting strategy to elicit comprehensive sentiment-related knowledge from the teacher LLM and systematically transfer this knowledge to the student model. The second stage, termed in-context learning distillation (ICLDist), transfers prompt-following capabilities in sentiment analysis to optimize the student model’s task alignment. In ICLDist, we enable the student model to follow task-specific instructions and demonstrations by mimicking the teacher LLM’s responses on few-shot samples. When constructing few-shot samples, we implement format and task diversification strategies to strengthen the generalization of ICLDist.
To facilitate a systematic evaluation, we develop SentiBench, a comprehensive sentiment analysis benchmark, comprising 3 task categories across 12 datasets. Our extensive experimentation on this benchmark reveals several key findings: (1) Our approach demonstrates substantial advantages over generic distillation methods, achieving effective distillation of LLMs’ sentiment analysis capabilities. Specifically, the student model achieves a 10% improvement in the average $F_{1}$-score across various tasks, with a particularly remarkable gain of 38% in irony detection. (2) Our approach enables the 1.2B model to outperform the original 3.2B model, and the 3.2B model to surpass the original 8B model. As illustrated in Figure 1, the resulting models exhibit strong competitiveness against other small-scale models. (3) Further analysis reveals the complementary nature of KnowDist and ICLDist and validates the effectiveness of each component in our approach.
@@ 2 Two-stage Distillation Framework
$\hat{\theta}_{S}$ $=\underset{\theta_{S}}{\operatorname{argmax}}\hskip-5.0pt\sum_{\bm{i},\bm{d},\bm{x},\bm{y}}\hskip-5.0pt\log P_{\cal M}(\bm{y}\ |\ \bm{i},\bm{d},\bm{x};\theta_{S}),$ (2)
[…]
@@ 2.2 In-Context Learning Distillation
[2pt/4pt] Aspect term sentiment analysis Rest16 1600 400 676 - micro_f1
@@
Structured sentiment analysis Opener 1744 249 499 - sentiment_graph_f1
Table 1: Task overview and dataset statistics in SentiBench. We perform downsampling on some datasets to ensure computational efficiency. For sampling details, please refer to Appendix B.1.
@@
The ICLDist process is illustrated in Figure 2. Similar to knowledge collection, we first gather a large volume of user-generated content. Next, we select sentiment classification and emotion recognition as distillation tasks [^3] For these two tasks, we construct a collection of 55 demos. Specifically, we first randomly select user-generated text of a suitable length, then use GPT-4o to annotate its sentiment and emotion labels, and finally perform manual verification to ensure the quality of the annotations. and construct prompts by randomly applying our format diversification strategies. Additionally, we incorporate the task diversification strategy to generate supplementary prompts. We then collect the teacher LLM’s responses to these prompts, resulting in a task-alignment corpus. Finally, we optimize the student model on this corpus to enhance its task alignment.
@@ 3 SentiBench
[2pt/4pt] Qwen-2.5-1.5B 91.92 97.30 92.33 52.39 65.80 63.61 70.90 35.73 37.66 53.25 18.47 20.08 58.29
 + Distill. w/ Alpaca-data 92.07 96.63 92.25 51.84 66.24 54.76 70.31 28.20 41.50 57.15 18.76 18.72 57.37(-0.92)
 + Distill. w/ Lamini-data 92.80 97.60 93.08 57.75 71.94 51.37 71.54 29.10 40.74 57.87 18.81 15.27 58.16(-0.13)
 + Know & ICLDist 93.80 98.10 95.99 65.89 71.69 72.57 74.83 47.31 51.36 53.98 22.93 19.70 64.01(+5.72)
[2pt/4pt] Llama-3-3.2B 92.57 96.53 93.59 61.45 64.00 68.88 71.43 33.32 46.37 51.66 11.09 23.10 59.50
 + Distill. w/ Alpaca-data 92.37 97.37 93.92 57.70 66.59 64.47 72.05 28.70 44.70 50.77 14.19 24.63 58.96(-0.54)
 + Distill. w/ Lamini-data 92.80 97.33 94.91 62.07 70.10 65.61 72.49 40.28 50.29 52.62 16.06 27.36 61.83(+2.33)
 + Know & ICLDist 94.30 98.17 95.41 69.57 85.25 77.47 75.10 48.24 53.24 66.01 22.95 35.16 68.41(+8.91)
Table 2: Experimental results on SentiBench ($F_{1}$-score, %).
@@ 4.1 Experimental Setup
Implementation Details. The teacher LLM is set to Llama-3.1-70B-Instruct Grattafiori et al. (2024), while Llama-3.2-1.2B-Instruct, Qwen-2.5-1.5B-Instruct [^4] https://qwenlm.github.io/blog/qwen2.5/, and Llama-3.2-3.2B-Instruct are employed as student models. For distillation, we curate a large and diverse corpus of user-generated texts from IMDb Nguyen et al. (2014), Yelp [^5] https://www.yelp.com/dataset, Amazon [^6] https://nijianmo.github.io/amazon/index.html, and Twitter [^7] https://archive.org/details/twitterstream. We preprocess this corpus by decontaminating it for the downstream datasets and eliminating duplicates using simhash. We then apply the proposed prompting methods to these user texts and obtain 1M KnowDist samples and 400K ICLDist samples. We further supplement the ICLDist corpus with 100K general task samples from the super-naturalInstruction Wang et al. (2022) dataset. The 1.2B and 1.5B models are optimized using the complete training set, while the 3.2B model is trained on a subset containing 200K KnowDist samples and 100K ICLDist samples. The hyperparameter settings are provided in Appendix C.1.
After distillation, we evaluate the student model on SentiBench using in-context learning, with dataset statistics shown in Table 1. The specific prompts for each task are detailed in Appendix B.2. During evaluation, we randomly sample 4 examples from the dev set as demonstrations. To ensure generation stability, we set the temperature parameter to 0 during model inference. To mitigate the impact of randomness, we conduct each evaluation using 3 different random seeds and report the average results.
Baselines. We compare our approach with generic distillation methods. Specifically, we train the student model using existing instruction-following datasets, including the 52K data constructed by Taori et al. (2023) (alpaca-data), and the 2.58M data developed by Wu et al. (2024) (lamini-data). Besides, we evaluate a diverse set of models for reference: (1) Llama-3 series models, spanning different scales (8B and 70B variants); (2) several small-scale models ranging from 1B to 3B parameters, including OPT-1.3B Zhang et al. (2022), TinyLlama-1.1B-Chat-v1.0 Zhang et al. (2024a), Phi-2-2.7B [^8] https://huggingface.co/microsoft/phi-2, and Gemma-2-2.6B-it Team (2024); and (3) GPT-3.5999Available at https://chat.openai.com/. The specific model used is gpt-3.5-turbo-0125..
@@ 4.2 Main Results
Table 3 presents the comparison results on SentiBench. We observe that two generic distillation methods yield only marginal and unstable gains in sentiment analysis performance, with the student model showing average $F_{1}$-score improvements below 2.33%. These limited improvements suggest that utilizing generic distillation methods to transfer sentiment analysis capabilities is ineffective. In contrast, our approach, namely Know & ICLDist, significantly enhances the sentiment analysis performance of the student model. Specifically, our approach achieves an average improvement of 10.33%, 5.72%, and 8.91%. The most striking improvement is observed in irony detection of Llama-3.2-1.2B-Instruct, where the $F_{1}$-score increases dramatically from 35.80% to 73.80% - an improvement of 38.00%. These results demonstrate the effectiveness of our approach in transferring sentiment analysis capabilities from the LLM to its more efficient counterparts.
Furthermore, the experimental results in Table 3 reveal several additional insights. Firstly, within the Llama-3 family, we observe a clear positive correlation between model size and performance, with Llama-3-70B achieving the best results, surpassing GPT-3.5. Secondly, our approach empowers the 1.2B model to outperform the original 3.2B model, and the 3.2B model to surpass the original 8B model. Moreover, the distilled models demonstrate strong competitive performance compared to other small-scale models and GPT-3.5, also illustrated in Figure 1. Thirdly, our approach demonstrates consistent performance gains across different model families, including Llama-3 and Qwen2.5, highlighting its broad generalization. Fourthly, with Llama-3-70B as the teacher LLM, our approach enables Llama-3-3.2B to achieve comparable performance to the teacher on sentiment classification, irony detection, and emotion recognition. Finally, both the distilled models and other small-scale models show inferior performance on intimacy analysis and tuple extraction tasks (i.e., ASQP and SSA). These tasks require a deep understanding of social context and advanced structured extraction capabilities, presenting promising directions for future research.
@@ 4.3 Analyis of Two-stage Optimization
Our distillation framework consists of two stages: KnowDist and ICLDist. Below, we conduct an in-depth analysis of these two stages, aiming to distinguish their respective roles and investigate how they complement each other.
@@
For the first question, we conduct fine-tuning experiments using the training samples from SentiBench. The results in Table 3 demonstrate that under the fine-tuning setting, both KnowDist and ICLDist can enhance the student model’s sentiment analysis performance. Notably, KnowDist achieves more substantial improvements, which contrasts with the in-context learning results in Figure 3. These findings support our claims: KnowDist strengthens the student model’s fundamental sentiment analysis capabilities, while ICLDist optimizes task alignment. When sufficient labeled samples are available for downstream task alignment, the benefits of ICLDist’s task alignment become less significant. However, such labeled data is often scarce in real-world applications. Consequently, both KnowDist and ICLDist are essential components of our framework.
@@ 4.4 Ablation Studies
Table 5: Ablation results of KnowDist ($F_{1}$-score, %). Anl and Rw denote analyzing and rewriting respectively, and MPP stands for multi-perspective prompting. The distillation samples used are 200K.
@@
Table 6: Ablation results of ICLDist ($F_{1}$-score, %). LW, LT, and MI denote the format diversification of Label Words, Label Taxonomies, and Minimized Instructions respectively, while TD represents Task Diversification. We divide tasks in SentiBench into seen and unseen categories during distillation, where seen tasks include sentiment classification and emotion recognition, while the rest are considered unseen. The distillation samples used are 100K.
@@ 4.5 Discussions
Effect of Teacher LLMs. We experiment with different teacher LLMs in our distillation framework to analyze their impact. The results in Table 7 reveal that teacher quality significantly influences distillation effectiveness, as larger teacher LLMs generally lead to more substantial improvements. Furthermore, we make two noteworthy observations. First, even when using identical models for both teacher and student, distillation has the potential to enhance the student’s sentiment analysis performance. This result suggests the potential for leveraging distillation to achieve self-improvement in specialized domains. Second, larger teachers do not always lead to better performance, as evidenced in FSA tasks, where the 8B teacher slightly outperforms the 70B teacher. We hypothesize that larger teachers may sometimes pose greater learning challenges for smaller student models, warranting further exploration in future work.
@@
Table 8: Experimental results on 5-shot MMLU (accuracy, %). Our evaluation is conducted using LM-Evaluation-Harness provided at https://github.com/meta-llama/llama-cookbook.